OpenStudy (anonymous):

Find the y-intercept of the function: f(x) = 6˟ + 2

jimthompson5910 (jim_thompson5910):

Plug in x = 0 and evaluate to find the y-intercept

OpenStudy (anonymous):

i still dont understand\

jimthompson5910 (jim_thompson5910):

f(x) = 6^x + 2 f(0) = 6^0 + 2 f(0) = ??

OpenStudy (anonymous):

i got 2 too but my teach has the choices written in pair

OpenStudy (anonymous):

teacher

jimthompson5910 (jim_thompson5910):

6^0 = ???

OpenStudy (anonymous):

0?

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

any number (but zero) to the zeroth power is always 1

jimthompson5910 (jim_thompson5910):

so 6^0 = 1

OpenStudy (anonymous):

oh so is the pair (0,1)

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

6^0 is just part of it

jimthompson5910 (jim_thompson5910):

remember the +2 at the end

OpenStudy (anonymous):

uggh im sorry this is just kinda confusing

jimthompson5910 (jim_thompson5910):

remember we had 6^0 + 2

jimthompson5910 (jim_thompson5910):

replace the 6^0 with 1 since the two are equal

jimthompson5910 (jim_thompson5910):

then evaluate

OpenStudy (anonymous):

8?

jimthompson5910 (jim_thompson5910):

6^0 = 1, so f(x) = 6^x + 2 f(0) = 6^0 + 2 f(0) = 1 + 2 I'm sure you see what to do from here

OpenStudy (anonymous):

f(0)=3

OpenStudy (anonymous):

(0,3)

jimthompson5910 (jim_thompson5910):

yep

jimthompson5910 (jim_thompson5910):

you got it

OpenStudy (anonymous):

thank you

jimthompson5910 (jim_thompson5910):

yw